OpenStudy (anonymous):

A Ferris wheel has a radius of 80 feet. Two particular cars are located such that the central angle between them is 165º. To the nearest tenth, what is the length of the intercepted arc between those two cars on the Ferris wheel?

OpenStudy (anonymous):

165/360 = 2.181818181818182 2 * 80 = 160 2.181818181818182/160 = 73.33333333333333 pi